If a boat goes downstream 72 miles in 3 hours and upstream 60 miles in 6 hours, the rate of the river and the rate of the boat in still water respectively are?
Answer by Alan3354(69443)

72/3 = 24 mph
60/6 = 10 mph
The boat's speed is the average = (24+10)/2 = 17 mph
The current is the difference = 24-17 = 7 mph